Consider an economy described by the following equations:
Y = C + I + G
Y = 5,000 G = 1,000 T = 1,000
C = 250 + 0.75(Y − T) I = 1,000 − 50 r.
a. In this economy, compute private saving, public saving, and national saving.
b. Find the equilibrium interest rate.
c. Now suppose that G rises to 1,250. Compute private saving, public saving, and national saving.
d. Find the new equilibrium interest rate.
1
Expert's answer
2020-12-08T07:32:44-0500

(a)private saving= Y-T-C

= 5000-1000-250-0.75(5000-1000) = 750

public saving = T-G

= 1000-1000 = 0

national saving = S = private saving+ public saving = 750


(B)

S+I

750= 1000-50r

r = 5%


(c) private saving is unchanged

public saving = 1000-1250 = - 250


(d) 500= 1000-50r

r = 10%